I'm doing elmination and I don't get how to do this problem I tried an keep getting te wrong answer 9x+5y=2 and -11x+2y=-27 can you please explain how to do it step by step please thanks
Found 2 solutions by mananth, Maths68:
Answer by mananth(16946)

9 x + 5 y = 2 .............1
-11 x + 2 y = -27 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 5
-18 x -10 y = -4
-55 x + 10 y = -135
Add the two equations
-73 x = -139
/ -73
x=1.9
plug value of x in (1)
9 x + 5 y = 2
17.14 + 5 y = 2
5 y = 2-17.14
5 y = -15.14
y = -3.03

Answer by Maths68(1474)

9x+5y=2 ……….(1)
-11x+2y=-27…………(2)
Multiply (1) by 2 and (2) by (-5) then add both equations.
18x+10=4
55x-10y=135
------------------
73x=139
x=139/73
Put the value of x in equation (1)
9(139/73)+5y=2
1251/73+5y=2
5y=2-1251/73
Taking LCM on Right hand side
5y=(146-1251)/73
y=-1105/73*5
Y=-1105/73*5
y=-221/73
x=139/73, y=-221/73
Check:
Put the values of x and y in equation (1)
9x+5y=2
9(139/73)+5(-221/73)=2
1251/73-1105/73=2
Multiply by 73 both sides
1251-1105=2*73
146=146
Now
Put the values of x and y in equation (2)
-11x+2y=-27
-11(139/73)+2(-221/73)=-27
-1529/73-442/73=-27
Multiply by 73 both sides
-1529-442=-27*73
-1971=-1971